If npgreater than or equals5 and nqgreater than or equals​5, estimate Upper P left parenthesis at least 5 right parenthesis with nequals13 and pequals0.6 by using the normal distribution as an approximation to the binomial​ distribution; if npless than5 or nqless than​5, then state that the normal approximation is not suitable. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.
Answer #1

Answer:

Given,

sample n = 13

p = 0.6

q = 1 - p

= 1 - 0.6

= 0.4

np = 13*0.6

= 7.8 > 5

nq = 13*0.4

= 5.2 > 5

The normal approximation is not suitable

Mean = np = 7.8

Standard deviation = sqrt(npq)

= sqrt(13*0.6*0.4)

= 1.77

P(X >= 4.5) = P((x-u)/s >= (4.5 - 7.8)/1.77)

= (z >= -1.86)

= 0.9685572 [since from z table]

= 0.9686